OpenStudy (anonymous):

Evaluate log4 16. −4 −2 2 4

OpenStudy (anonymous):

I want to say its 2 because 4^2 = 16

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

log(16)/ log(4) = 2

OpenStudy (anonymous):

2^4 = 16 as a double check

OpenStudy (anonymous):

er yah the other way around. the way u did it is right

OpenStudy (anonymous):

let \[\log_{4} 16=x,4^{x}=16=4^{2},x=2\]

OpenStudy (whpalmer4):

Yes, \[\log_4 16=2\]because \[4^2 = 16\]

OpenStudy (anonymous):

sweet, I have two more questions if your up for them?